OpenStudy (anonymous):

use the fundamental identities to solve the equation. 1/tan^2x+1

OpenStudy (valpey):

That isn't really an equation. Equations tend to have equals signs.

OpenStudy (valpey):

Also do you mean: \[\frac{1}{tan^2(x+1)}\] \[\frac{1}{tan^2(x)+1}\] \[\frac{1}{tan^2(x)}+1\] Or something else?

OpenStudy (anonymous):

oops sorry!! simplify the expression*